an atom emits a spectral line of wavelength lambda when an electron makes a transition between levels of energy E1 and E2. Which expression correctly relates lamba, E1 and E2?
 

Grade:12th Pass

1 Answers

Vikas TU
4 years ago
Dear student 
We knw that 
E = hc / lambda 
Energy difference = E1-E2 
E1 = hc / lambda 1 
E2 = hc / lambda 2 
E1 – E2 = hc (1/lambda1 – 1/lambda2 )
Hope this helps
Good luck